An acquirer with a P/E ratio of 13 and earnings of $1.95 seeks to take over another target firm with value $17.6 and P/E ratio 16. What is the new merged firm's P/E?

Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Value of the acquirer = P/E*earnings

= 13*1.95

= $25.35

Now earnings of target firm = Value/(P/E) = 17.6/16 = 1.10

So, combined earnings = 1.95+1.10 = 3.05

And combined value = 25.35+17.6 = 42.95

Thus merged firm's P/E = combined value/combined earnings

= 42.95/3.05

= 14.0820